# Spectral finiteness, quantum norm continuity and classical points

### 12 Mar 2026 v1

@@ Introduction
There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1. [8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-5 below). In that Banach-space context, said norm (or uniform) continuity is defined via [8, Lemma 3.7] as
@@ 1 Quantum uniform continuity in several guises, and consequences
(1-1) $\Braket{x\mid y}_{r}:=h(x^{*}y)\quad\text{and}\quad\Braket{x\mid y}_{\ell}:=h(xy^{*}),$
@@
(1-2) $x^{\alpha}_{ij}:=\frac{q_{\alpha}}{F^{-1}_{\alpha ii}}(h\otimes\operatorname{id})\left(u^{\alpha*}_{ij}x\right)=\frac{q_{\alpha}}{F_{\alpha jj}}\left(h\otimes\operatorname{id})(xu^{\alpha*}_{ij}\right),$
@@
(1-3) $E\left(xx^{*}\right):=\left(h\otimes\operatorname{id}\right)(xx^{*})$ $=\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}h\left(u^{\alpha}_{ij}u^{\beta*}_{kl}\right)x^{\alpha}_{ij}x^{\beta*}_{kl}$
@@
It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to. Reprising the notation of ˜1-3 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side. Applied to $x:=U$, ˜1-3 reads
@@
(1-4) $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})U^{*}\in{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}})\ni u^{\beta}_{ij}\xmapsto{\quad\psi^{\alpha}\quad}\delta_{\alpha\beta}\delta_{ii}\in{\mathbb{C}}$
@@
  • ${\mathcal{C}}({\mathbb{G}})\xrightarrow{\pi_{h}}{\mathcal{L}}\left(L^{2}({\mathbb{G}})\right)$ for the GNS representation of the Haar state $h\in{\mathcal{C}}({\mathbb{G}})^{*}$: the completion of ${\mathcal{O}}({\mathbb{G}})$ under the inner product $\Braket{-\mid-}_{r}$ of ˜1-1;
@@
One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-4 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
(1-6) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
@@
A Banach-space representation ˜1-5 of a compact quantum group is boundedly uniform or uniform≤1 if the continuity in ˜0-1 holds on norm-bounded subsets of ${\mathcal{C}}({\mathbb{G}})^{*}$ (equivalently, on any one norm-bounded origin neighborhood, e.g. the unit ball). $\blacklozenge$
@@
For the latter map this is immediate given the conversion procedure ˜1-6 from unitary to Banach-space representations, while the passage between the two is effected by $\left(\left(\psi\otimes\operatorname{id}\right)U\right)^{*}=\left(\psi^{*}\otimes\operatorname{id}\right)U^{*}$ for $\psi^{*}:=\overline{\psi\left(\bullet^{*}\right)}$. $\blacksquare$
@@
(1-7) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
@@
  1. ˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-2 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-1, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal). The hypothesis ensures the existence of norm-1
@@
  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ is
@@
  • The terminology applies here (somewhat unconventionally) also “locally” or relatively, to families of representations: given any finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ one might consider sequences $\alpha_{n}$ exhibiting rapid decay with respect to $\ell:=\ell_{{\mathcal{F}}}$ in a sense the reader will have no difficulties extrapolating from the global version just sketched.
@@
Let $\mathbb{G}$} be a compact quantum group, ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ a finite family and $\left(\alpha_{n}\right)_{n}\subset\operatorname{Irr}({\mathbb{G}})$ with property RD with respect to $\ell_{{\mathcal{F}}}$ and trivial $F^{\alpha_{n}}$.
@@
Write $H:=\bigoplus_{n}H^{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
(1-8) $\|v\|=1\ \wedge\ \operatorname{\mathrm{supp}}v\subseteq\left\{\alpha_{m}\ :\ m\gg 0\right\}$
($\operatorname{\mathrm{supp}}$ defined as in ˜1-7) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large. The unimodularity constraint $F^{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
˜1-8 $\xRightarrow{\quad}\ \exists\left(c>1\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|_{2}\leq\max_{m:\alpha_{m}\in\operatorname{\mathrm{supp}}v}\frac{1}{\sqrt{\dim\alpha_{m}}}\leq\frac{1}{c^{m_{0}}}\right)$

### 13 Mar 2026 v2

@@ Introduction
There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1. [8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-6 below). In that Banach-space context, said norm (or uniform) continuity is defined via [8, Lemma 3.7] as
@@ 1 Quantum uniform continuity in several guises, and consequences
(1-2) $\Braket{x\mid y}_{r}:=h(x^{*}y)\quad\text{and}\quad\Braket{x\mid y}_{\ell}:=h(xy^{*}),$
@@
(1-3) $x^{\alpha}_{ij}:=\frac{q_{\alpha}}{F^{-1}_{\alpha ii}}(h\otimes\operatorname{id})\left(u^{\alpha*}_{ij}x\right)=\frac{q_{\alpha}}{F_{\alpha jj}}\left(h\otimes\operatorname{id})(xu^{\alpha*}_{ij}\right),$
@@
(1-4) $E\left(xx^{*}\right):=\left(h\otimes\operatorname{id}\right)(xx^{*})$ $=\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}h\left(u^{\alpha}_{ij}u^{\beta*}_{kl}\right)x^{\alpha}_{ij}x^{\beta*}_{kl}$
@@
It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to. Reprising the notation of ˜1-4 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side. Applied to $x:=U$, ˜1-4 reads
@@
(1-5) $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})U^{*}\in{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}})\ni u^{\beta}_{ij}\xmapsto{\quad\psi^{\alpha}\quad}\delta_{\alpha\beta}\delta_{ii}\in{\mathbb{C}}$
@@
  • ${\mathcal{C}}({\mathbb{G}})\xrightarrow{\pi_{h}}{\mathcal{L}}\left(L^{2}({\mathbb{G}})\right)$ for the GNS representation of the Haar state $h\in{\mathcal{C}}({\mathbb{G}})^{*}$: the completion of ${\mathcal{O}}({\mathbb{G}})$ under the inner product $\Braket{-\mid-}_{r}$ of ˜1-2;
@@
One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-5 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
(1-7) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
@@
A Banach-space representation ˜1-6 of a compact quantum group is boundedly uniform or uniform≤1 if the continuity in ˜0-1 holds on norm-bounded subsets of ${\mathcal{C}}({\mathbb{G}})^{*}$ (equivalently, on any one norm-bounded origin neighborhood, e.g. the unit ball). $\blacklozenge$
@@
For the latter map this is immediate given the conversion procedure ˜1-7 from unitary to Banach-space representations, while the passage between the two is effected by $\left(\left(\psi\otimes\operatorname{id}\right)U\right)^{*}=\left(\psi^{*}\otimes\operatorname{id}\right)U^{*}$ for $\psi^{*}:=\overline{\psi\left(\bullet^{*}\right)}$. $\blacksquare$
@@
(1-8) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
@@
  1. ˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-3 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-2, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal). The hypothesis ensures the existence of norm-1
@@
  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ (typically assumed symmetric, i.e. with the dual $\alpha^{*}\in{\mathcal{F}}$ whenever $\alpha\in{\mathcal{F}}$) is
@@
  • The terminology applies here (somewhat unconventionally) also “locally” or relatively, to families of representations: given any symmetric finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ one might consider sequences $\alpha_{n}$ exhibiting rapid decay with respect to $\ell:=\ell_{{\mathcal{F}}}$ in a sense the reader will have no difficulties extrapolating from the global version just sketched.
@@
Let $\mathbb{G}$} be a compact quantum group, ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ a finite family and $\left(\alpha_{n}\right)_{n}\subset\operatorname{Irr}({\mathbb{G}})$ with property RD with respect to $\ell_{{\mathcal{F}}}$.
@@
Observe first that the individual positive operators $F_{\alpha_{n}}$ of ˜1-1 are identities: [31, Proposition 4.7] proves this globally, assuming the quantum group as a whole has property RD, but the argument (relying on [31, Lemmas 4.3 and 4.6]) goes through in the local version needed here.
Write $H:=\bigoplus_{n}V_{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
(1-9) $\|v\|=1\ \wedge\ \operatorname{\mathrm{supp}}v\subseteq\left\{\alpha_{m}\ :\ m\gg 0\right\}$
($\operatorname{\mathrm{supp}}$ defined as in ˜1-8) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large. The unimodularity condition $F_{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
˜1-9 $\xRightarrow{\quad}\ \exists\left(c>1\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|_{2}\leq\max_{m:\alpha_{m}\in\operatorname{\mathrm{supp}}v}\frac{1}{\sqrt{\dim\alpha_{m}}}\leq\frac{1}{c^{m_{0}}}\right)$